OpenStudy (anonymous):

An electron is trapped in a box of length L = 100 pm. There is a single node in the center of the box where the electron cannot exist. What is the energy of the electron, in eV?

OpenStudy (frostbite):

For a one dimensional particle in a box problem, the total solution for the Schrödinger equation's energy is \(\Large E_{n}=\frac{n^{2}h^2}{8mL^2}, n=1,2,3....\)

OpenStudy (frostbite):

Solve the problem using the correct wavefuction ( the correct \(n\))

OpenStudy (aaronq):

Frosty to the rescue! though, i think you mean principal quantum number, not wavefunction at the end there :P

OpenStudy (anonymous):

please solve it... I am beginner

OpenStudy (frostbite):

No actually I meant wavefunction :P remember that n is am integer we put on our waveterm in order to hit all the allowed solutions. (see figure). \[\Large n \times \frac{ 1 }{ 2 } \lambda = L, ~ n=1,2,3....\]